The assembly has the diameters and material make-up indicated. It fits securely between its fixed supports when the temperature is T1 = 70?F,determine the average normal stress in each material when the temperature reaches T2 = 108
Answer #1
Concepts and reason

The concepts required to solve the question are stress, strain, and thermal expansion.

First, write the equilibrium condition of the deflection of the bar by using the expression of the deflections due to stress and strain, and thermal expansion to find the force due to fixed ends.

Use the calculated force to find the normal stress in each bar of the material.

Stress: It is defined as the internal force which is offered by a material per unit area.

Strain: It is defined as the change in length of material to the original length of material.

Hooke’s law: Within elastic limit, stress is directly proportional to strain.

Thermal expansion: In thermal expansion, material changes its shape and size due to change in temperature.

Fundamentals

Write the expression of the stress.

σ=PA\sigma = \frac{P}{A}

Here, σ\sigma is the stress, PP is the internal force offered by material and AA is the area of material at which force acts.

Write the expression of the strain.

ε=δL\varepsilon = \frac{\delta }{L}

Here, ε\varepsilon is the strain, δ\delta is the change in length and LL is the original length.

From Hooke’s law, the stress can be written as,

σεσ=EεE=σε\begin{array}{c}\\\sigma \propto \varepsilon \\\\\sigma = E\varepsilon \\\\E = \frac{\sigma }{\varepsilon }\\\end{array}

Here, EE is the modulus of elasticity.

Substitute PA\frac{P}{A} for σ\sigma and δL\frac{\delta }{L} for ε\varepsilon .

E=(PA)(δL)E=PLδAδ=PLEA\begin{array}{c}\\E = \frac{{\left( {\frac{P}{A}} \right)}}{{\left( {\frac{\delta }{L}} \right)}}\\\\E = \frac{{PL}}{{\delta A}}\\\\\delta = \frac{{PL}}{{EA}}\\\end{array}

Write the expression of change in length due to thermal expansion.

δt=αL(T2T1){\delta _t} = \alpha L\left( {{T_2} - {T_1}} \right)

Here, δt{\delta _t} is the change in length due to thermal expansion, α\alpha is the coefficient of thermal expansion, LL is the length of material, T1{T_1} is the initial temperature and T2{T_2} is the final temperature.

Write the expression of the cross-sectional area of the bar.

A=πd24A = \frac{{\pi {d^2}}}{4}

Here, AA is the area of the bar and dd is the diameter of the bar.

(A)

Free body diagram of the assembly can be shown as,

2014-16
Aluminum
304 Stainless
Steel
C 86100
Bronze
F-
12 in
8in
4 in
F
Kan * 67 +387

From above diagram, the displacement between AA to DD is zero because both the ends are fixed that can be written as,

δAD=0{FLABAABEAl+αAlLAB(T2T1)+FLBCABCEBr+αBrLBC(T2T1)+FLCDACDESt+αStLCD(T2T1)=0}{4FLABπdAB2EAl+αAlLAB(T2T1)4FLBCπdBC2EBr+αBrLBC(T2T1)4FLCDπdCD2ESt+αStLCD(T2T1)=0}\begin{array}{l}\\{\delta _{AD}} = 0\\\\\left\{ \begin{array}{l}\\\frac{{ - F{L_{AB}}}}{{{A_{AB}}{E_{Al}}}} + {\alpha _{Al}}{L_{AB}}\left( {{T_2} - {T_1}} \right) + \frac{{ - F{L_{BC}}}}{{{A_{BC}}{E_{Br}}}}\\\\ + {\alpha _{Br}}{L_{BC}}\left( {{T_2} - {T_1}} \right) + \frac{{ - F{L_{CD}}}}{{{A_{CD}}{E_{St}}}} + {\alpha _{St}}{L_{CD}}\left( {{T_2} - {T_1}} \right) = 0\\\end{array} \right\}\\\\\left\{ \begin{array}{l}\\ - \frac{{4F{L_{AB}}}}{{\pi d_{AB}^2{E_{Al}}}} + {\alpha _{Al}}{L_{AB}}\left( {{T_2} - {T_1}} \right) - \frac{{4F{L_{BC}}}}{{\pi d_{BC}^2{E_{Br}}}}\\\\ + {\alpha _{Br}}{L_{BC}}\left( {{T_2} - {T_1}} \right) - \frac{{4F{L_{CD}}}}{{\pi d_{CD}^2{E_{St}}}} + {\alpha _{St}}{L_{CD}}\left( {{T_2} - {T_1}} \right) = 0\\\end{array} \right\}\\\end{array}

Here, FF is the force that applied on the material, LAB{L_{AB}} is the length of the aluminum bar, AAB{A_{AB}} is the area of the bar ABAB , EAl{E_{Al}} is the modulus of elasticity of aluminium bar, αAl{\alpha _{Al}} is the coefficient of the thermal expansion, dAB{d_{AB}} is the diameter of the aluminum bar and similarly for bronze and stainless steel.

From the experimental data, the coefficient of thermal expansion of 2014T6{\rm{2014 - T6}} aluminium is 12.8×106/F12.8 \times {10^{ - 6}}{\rm{/^\circ F}} , the coefficient of thermal expansion of C86100{\rm{C86100}} bronze is 9.60×106/F9.60 \times {10^{ - 6}}{\rm{/^\circ F}} , and the coefficient of thermal expansion of 304304 stainless steel is 9.6×106/F9.6 \times {10^{ - 6}}{\rm{/^\circ F}} .

From the experimental data, the modulus of elasticity of the aluminum material is 10.6×106psi10.6 \times {10^6}{\rm{ psi}} , the modulus of elasticity of the bronze material is 15×106psi15 \times {10^6}{\rm{ psi}} , and the modulus of elasticity of stainless steel material is 28×106psi28 \times {10^6}{\rm{ psi}} .

Substitute (4×12)in\left( {4 \times 12} \right){\rm{ in}} for LAB{L_{AB}} , 12in12{\rm{ in}} for dAB{d_{AB}} , 10.6×106psi10.6 \times {10^6}{\rm{ psi}} for EAl{E_{Al}} , 12.8×106/F12.8 \times {10^{ - 6}}{\rm{/^\circ F}} for αAl{\alpha _{Al}} , (6×12)in\left( {6 \times 12} \right){\rm{ in}} for LBC{L_{BC}} , 8in{\rm{8 in}} for dBC{d_{BC}} , 15×106psi15 \times {10^6}{\rm{ psi}} for EBr{E_{Br}} , 9.60×106/F9.60 \times {10^{ - 6}}{\rm{/^\circ F}} for αBr{\alpha _{Br}} , (3×12)in\left( {3 \times 12} \right){\rm{ in}} for LCD{L_{CD}} , 4in{\rm{4 in}} for dCD{d_{CD}} , 28×106psi28 \times {10^6}{\rm{ psi}} for ESt{E_{St}} , 9.60×106/F9.60 \times {10^{ - 6}}{\rm{/^\circ F}} for αSt{\alpha _{St}} , 70F70{\rm{^\circ F}} for T1{T_1} and 108F108{\rm{^\circ F}} for T2{T_2} . [4((4×12)in)π(12in)2(10.6×106psi)+(12.8×106/F)((4×12)in)(108F70F)4F((6×12)in)π(8in)2(15×106psi)+(9.60×106/F)((6×12)in)(108F70F)4F((3×12)in)π(4in)2(28×106psi)+(9.60×106/F)((3×12)in)(108F70F)]=0[4×108F+23347.2×1069.55×108F+26265.6×10610.23×108F+13132.8×106]=0\begin{array}{c}\\\left[ \begin{array}{l}\\ - \frac{{4\left( {\left( {4 \times 12} \right){\rm{ in}}} \right)}}{{\pi {{\left( {12{\rm{ in}}} \right)}^2}\left( {10.6 \times {{10}^6}{\rm{ psi}}} \right)}} + \left( {12.8 \times {{10}^{ - 6}}{\rm{/^\circ F}}} \right)\left( {\left( {4 \times 12} \right){\rm{ in}}} \right)\left( {108{\rm{^\circ F}} - 70{\rm{^\circ F}}} \right)\\\\ - \frac{{4F\left( {\left( {6 \times 12} \right){\rm{ in}}} \right)}}{{\pi {{\left( {{\rm{8 in}}} \right)}^2}\left( {15 \times {{10}^6}{\rm{ psi}}} \right)}} + \left( {9.60 \times {{10}^{ - 6}}{\rm{/^\circ F}}} \right)\left( {\left( {6 \times 12} \right){\rm{ in}}} \right)\left( {108{\rm{^\circ F}} - 70{\rm{^\circ F}}} \right) - \\\\\frac{{4F\left( {\left( {3 \times 12} \right){\rm{ in}}} \right)}}{{\pi {{\left( {{\rm{4 in}}} \right)}^2}\left( {28 \times {{10}^6}{\rm{ psi}}} \right)}} + \left( {9.60 \times {{10}^{ - 6}}{\rm{/^\circ F}}} \right)\left( {\left( {3 \times 12} \right){\rm{ in}}} \right)\left( {108{\rm{^\circ F}} - 70{\rm{^\circ F}}} \right)\\\end{array} \right] = 0\\\\\left[ \begin{array}{l}\\ - 4 \times {10^{ - 8}}F + 23347.2 \times {10^{ - 6}}\\\\ - 9.55 \times {10^{ - 8}}F + 26265.6 \times {10^{ - 6}}\\\\ - 10.23 \times {10^{ - 8}}F + 13132.8 \times {10^{ - 6}}\\\end{array} \right]{\rm{ = 0}}\\\end{array}

Further, solve,

23.78×108F=62745.6×106F=263858.70lbf(1kip1000lbf)=263.8587kip\begin{array}{c}\\23.78 \times {10^{ - 8}}F = 62745.6 \times {10^{ - 6}}\\\\F = 263858.70{\rm{ lbf}}\left( {\frac{{1{\rm{ kip}}}}{{1000{\rm{ lbf}}}}} \right)\\\\ = 263.8587{\rm{ kip}}\\\end{array}

Calculate the average normal stress in the aluminum bar.

σAB=4FπdAB2{\sigma _{AB}} = \frac{{4F}}{{\pi d_{AB}^2}}

Substitute 12in12{\rm{ in}} for dAB{d_{AB}} and 263.8587kip263.8587{\rm{ kip}} for FF .

σAB=4(263.8587kip)π(12in)2=2.34ksi\begin{array}{c}\\{\sigma _{AB}} = \frac{{4\left( {263.8587{\rm{ kip}}} \right)}}{{\pi {{\left( {12{\rm{ in}}} \right)}^2}}}\\\\ = 2.34{\rm{ ksi}}\\\end{array}

(B)

Calculate the average normal stress in the bronze bar.

σBC=4FπdBC2{\sigma _{BC}} = \frac{{4F}}{{\pi d_{BC}^2}}

Substitute 8in{\rm{8 in}} for dBC{d_{BC}} and 263.8587kip263.8587{\rm{ kip}} for FF .

σBC=4(263.8587kip)π(8in)2=5.25ksi\begin{array}{c}\\{\sigma _{BC}} = \frac{{4\left( {263.8587{\rm{ kip}}} \right)}}{{\pi {{\left( {{\rm{8 in}}} \right)}^2}}}\\\\ = 5.25{\rm{ ksi}}\\\end{array}

Ans: Part A

The average normal stress in the aluminum bar is 2.34ksi2.34{\rm{ ksi}} .

Part B

The average normal stress in the bronze bar is 5.25ksi5.25{\rm{ ksi}} .